Shift report 20011025 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Thu Oct 25 2001 - 08:00:35 EDT

  • Next message: Djamel Ouerdane: "Re: TOFW Pid"

    ------------------------------------------------------------
    Date: Thu Oct 25 08:00:35 2001, Shift: 00:00-08:00
    Supervisor: CEJ, with: 
    Log book#10 pages 55 - 56
    
    Shift Summary:
    ------------------------------------------------------------
    1. Beam untill 2:20. Beam again from 3:20. BRAHMS ZDC rate
       is approx half of the other experiments.
    
    2. Runs:
       5509 :  78K tr-6 (continued from last shift)
       5512 : 100K tr-6
       5513 :  93K tr-6 (OBS: fs at 3 deg)
    
    3. All T5 anodes tripped in beginning of run5511. Turned 
       them on and started a new run. No other problems. 
    
    (Since there's beam study for the next 8 hours we've turned 
    of the magnets and left the building)
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Thu Oct 25 2001 - 08:01:25 EDT